Please help! I have attempted this question 9 times and I do not know what I am doing wrong. Thank You!

A projectile is fired at v0 = 368.0 m/s at an angle of θ = 63.0o with respect to the horizontal. Assume that air friction will shorten the range by 31.1%. How far will the projectile travel in the horizontal direction, R?
Answer #1

Initial Range, R' = u^2 sin(2 theta) /g

R' = 368^2 sin(2 x 63)/9.8 = 11179.62 m

Range, R = (1 - 0.311) x 11179.62 = 7702.76 m

Comment in case any doubt please rate my answer...